OpenStudy (anonymous):

Find the equation of the line that contains the point (27,-11) and is parallel to the line 2x+9y=18. Write your answer in slope intercept form. HOW DO YOU DO THIS!!!!

OpenStudy (amistre64):

yay!!!

OpenStudy (amistre64):

find the slope of the line and flip it over spank its bottom throw it down and but a negative to it

OpenStudy (dumbcow):

haha first find slope of line 2x+9y =18 get y by itself

OpenStudy (amistre64):

-a/b is the slope of yer line ;)

OpenStudy (dumbcow):

hold up amistre thats for perpendicular lines, this is a parallel line

OpenStudy (amistre64):

ack!! they need different words for those things; those are too similar lol

OpenStudy (amistre64):

in that case; erase the 18, fill in yer x and y and solve; then git rid of the x and y values

OpenStudy (amistre64):

2x+9y=18 ; chuck the 18...its useless to us 2x+9y= ; now fill in your (x=27,y=-11) and solve 2(27)+9(-11)=

OpenStudy (amistre64):

54 -99 = -45 2x+9y=-45

OpenStudy (dumbcow):

now put it in slope-intercept form by moving 2x over and dividing by 9 9y = -2x -45 y = (-2/9)x -5 that is the equation of the line going through point (27,-11) and parallel to given line

OpenStudy (anonymous):

thank you everyone, i really appreciate, will have a few more questions later on 2nite. Calc final is tomorrow @ 9am